Assume that a 2 kg of a certain gas with R=0.4305 kJ/kg-K undergoes a process that results in these changes: ΔU=1458.01 kJ, ΔH=2189.13 kJ. What is the value of cv?
1
Expert's answer
2021-07-12T03:36:42-0400

"\triangle H= C_P\triangle T"


"\triangle U= C_V\triangle T"


Specific heat ratio, "k = \frac{C_p}{C_v} =\frac {\triangle H}{\triangle U}= \frac {2189.13}{1458.01}=1.49"


"C_p-C_v = R" ​...(1)


"k = \frac{C_p}{C_v}" ​...(2)


From 1 and 2

"R = 1.49C_v-C_v = 0.49C_v"


"C_v = \frac{0.4305}{0.49}=0.8785" "\frac {kJ} {kg K}"